solve the inequality.
-18x-24≤-6(2x+1)

Answer by checkley77(12844)

-18x-24≤-6(2x+1)
-24<=18x-12x-6
-24+6<=6x
-18<=6x
-18/6<=x
-3<=x
Proof:
-18(-3)-24<=-6(2(-3)+1)
54-24<=-6(-6+1)
30<=-6*-5
30<=30